Determine the net force acting on the 50kg mass and express it in polar coordinates.
x= -12 m/s2 y= 8m/s2
1
Expert's answer
2018-11-12T15:56:09-0500

F_x=-12(50)=-600 N.

F_y=8(50)=400 N

In polar coordinates:

F=√(〖400〗^2+(-600)^2 )=721 N.

θ=180-tan^(-1)⁡〖400/600〗=146°